There was a problem loading the comments.

Etwatwa Zone 1,  Etwatwa Zone 2 INC-000024282

Support Portal  »  Fibre  »  Viewing Article

  Print

Vumatel is currently experiencing the following issue:

 

Start Date & Time: 26/09/2025 @ 15:03

Affected Area/s: Etwatwa Zone 1,  Etwatwa Zone 2  

Impact: Customers may experience no connectivity.

 

Update (03/10/2025 08:52):

Vumatel confirmed that this outage was resolved on 02/10/2025 22:12. Please restart your devices and contact our Support Team should you still experience connection issues.

 

Update (02/10/2025 @ 21:25):
Vumatel confirmed that this outage was resolved on 02/10/2025 @ 21:25. Please restart your devices and contact our Support Team at 010 593 2665 or support@coolideas.co.za should you still experience connection issues.

 

Update (02/10/2025 @ 20:04):
Technicians are actively addressing the remaining cases and making good headway. We’ll provide the next update at 22h00  . More updates to follow.
 

Update (02/10/2025 @16:05):

 Technicians are actively addressing the remaining cases and making good headway. We’ll provide the next update at 20h00 .

 

Update (02/10/2025 @13:18):

Technicians are making good progress with the cases that remain down. Next update will be provide at 16:00 

 

Update (02/10/2025 @ 12:30):

Remaining down services are being attended to.More updates to follow.

 

Update (02/10/2025 @ 12:27):

Associated tickets are being tested to ensure no further work is required. More updates to follow.

 

Update (02/10/2025 @ 08:43):

 Technicians are continuing to work on the remaining down cases this morning. The next update will be provided at 12:00.  . More updates to follow.

 

Update (01/10/2025 @ 22:29):

 The incident has been placed on hold. The team will continue in the morning at 8 am with repairs  . More updates to follow.

 

Update (01/10/2025 @ 19:06):

 Technicians continue to work on the outstanding fallout cases. The next update will be shared by 22h00  

 

Update (01/10/2025 @ 15:57):

 Technicians continue to work on the outstanding fallout cases. The next update will be shared by 19h00  . More updates to follow.

 

Update (01/10/2025 @12:10):

Technicians are actively working on the remaining fallout cases, further updates will be provided at 16h00  

 

Update (01/10/2025 @ 09:09):

Technicians are attending to the services that are down.Next update: 12h00  . More updates to follow.

 

Update (01/10/2025 @ 08:10):

 Technicians are enroute to site to continue with repairs.  . More updates to follow.

 

Update (01/10/2025 @ 00:42):

 Repair on hold due to area not safe technicians will attend the remaining fallouts in the morning . More updates to follow.

 

Update (30/09/2025 @ 20:37):

 Teams have completed with repairs and are now working on the services that remain down. Further updates will be provided at 22h30.  . More updates to follow.

 

Update (30/09/2025 @19:23):

 Repairs in progress - Splicing in progress on the 288f .  

 

Update (30/09/2025 @18:13):

 Teams are busy with floating in progress on the 48f backhaul link, 288f splicing on tube 7.  

 

Update (30/09/2025 @16:11):

 Teams are currently splicing tube 5 on the 288f. 72F splicing is complete.  

 

Update (30/09/2025 @16:11):

 Team is currently splicing - on Tube 2 on the 288f, and the 4th tube on the 72f.  

 

Update (30/09/2025 @14:59):

Site A: Team is currently still preparing and sizing 288f, 72f splicing has started,
Site B: 288f splicing has been completed  

 

Update (30/09/2025 @13:31):

Site A: Team is currently preparing 288f, 72f & 48f for splicing 
Site B: 288f splicing by Core AG completed  

 

Update (30/09/2025 @12:11):

Team is done trenching, exposing ducts and planting manhole, currently sizing slack to prepare cables for splicing  

 

Update (30/09/2025 @ 10:49):

 Team is pulling slack & planting manhole  . More updates to follow.

 

Update (30/09/2025 @ 09:52):

 Site A: Team is trenching to plant Rhinode 600
Site B: Finishing splicing from core ag side
 . More updates to follow.

 

Update (30/09/2025 @ 08:21):

 Technicians are on site localising the fault . More updates to follow.

 

Update (29/09/2025 @ 19:19):

 NWI is currently on hold. During cable floating, the duct was stacked at several points. The team will resume tomorrow with trenching and continue floating the cables to Point B in preparation for splicing.  

 

Update (29/09/2025 @18:04):

 Splicing still in progress on site A on both cables, 288f link & 72f link cables on 10th tray. Civil in progress on trenching to duct Exposed..  

 

Update (29/09/2025 @16:49):

 Splicing still in progress on Site A, 8th tray on 288f , team Trenching to unblock the duct for float the cables still stacked.

 

Update (29/09/2025 @15:22):

 Still splicing site A where the Cables start, we still have to float & pass where the incident was with all the cables, at the moment we are installing duct preparing for floating cable to pass to B site 48f core, 72f core and 288f link cable.  

 

Update (29/09/2025 @14:35):

Splicing still in progress on one site A, on 7 tube on 288f & 72f links. 
We still have to couple the 7way duct to prepare for site B.  

 

Update (29/09/2025 @13:08):

Splicing in progress: Splicing on tube 5 on 288f link cable. 
Auditing done on 72f.  

 

Update (29/09/2025 @11:20):

Splicing in progress: Splicing in progress on tube3 288f link cable.
Auditing in progress on 72f core cable.  

 

Update (29/09/2025 @ 10:25):

 Technicians onsite and assessing  . More updates to follow.

 

Update (29/09/2025 @ 08:38):

 Technicians are assessing the situation with regards to the weather conditions and continuing with repairs. As soon as further feedback has been received, an update will be provided.  . More updates to follow.

 

Update (28/09/2025 @ 18:49):

 Repairs on hold due to adverse weather conditions . More updates to follow.

 

Update (28/09/2025 @ 18:14):

 Technicians are on site preparing for repairs . More updates to follow.

 

Update (27/09/2025 @14:15):

The incident is currently on hold. 
Technicians are unable to proceed as clearance has not yet been granted by the TLB and municipal authorities. However, work will resume as soon as the necessary permissions are obtained.  

 

Update (27/09/2025 @13:16):

Stock has arrived on site. Civils is preparing to float the cable.  

 

Update (27/09/2025 @12:13):

Teams are waiting for stock to arrive so they can start with repair.  

 

Update (27/09/2025 @11:35):

Technicians are pulling slack  

 

Update (27/09/2025 @10:10):

Technicians are on site localising the fault  

 

Update (26/09/2025 @ 17:32):

The incident is currently on hold. The TLB and municipal authorities are not allowing the technicians to proceed, as the area needs to be cleared of water and the damaged pipe must be repaired before any further work can take place.

There is still no ETR at the moment.

 

Update: 26/09/2025 16:06

 Technicians are in the process of exposing the ducts to carry out repairs  


#Ekurhuleni #EtwatwaZone1 #Vumatel Reach


Share via

Related Articles

© Cool Ideas